# Phase-shifted remote photoplethysmography for estimating heart rate and blood pressure from facial video

### 09 Jan 2024 v1

@@ 4.1 Datasets
To demonstrate the heart rate and blood pressure estimation accuracy of the proposed method, we experimented on the MMSE-HR (Multimodal Spontaneous Expression-Heart Rate) (Zhang et al., 2016b) and V4V (Vision for Vitals) database (Revanur et al., 2021). The two datasets are sub-datasets of the MMSE database (BP4D+), which consists of synchronized facial image sequences and continuous blood pressure signals. As introduced in Section 3.1, the biometric information of PPG, heart rate, SBP, and DBP can be calculated by preprocessing the blood pressure signals. To ensure data diversity, tasks were performed to arouse amusement, pain, anger, and disgust. RGB video was recorded at the resolution of 1040$\times$1392 and the frame rate of 25 FPS, and blood pressure signals were collected at the sampling rate of 1000 Hz. The MMSE-HR database released at the CVPR 2016 for The Heart Rate Estimation Algorithm Challenge includes 102 samples from 17 male and 23 female subjects. The average signal lengths are 30 seconds. The V4V database released at the ICCV workshop 2021 for The 1st Vision for Vitals Challenge includes 1,358 data samples split into 724 training, 276 validation, and 358 test samples. The average signal lengths are 40 seconds, collected from 179 subjects.
@@
Figure 5: Pearson correlation graph of predicted and actual HR
Figure 6: Predicted facial(blue), acral(green) rPPG and PPG(red)
@@ 4.2 Quantitative results on MMSE-HR database
In this section, we compare the heart rate estimation performance of the proposed algorithm with previous methods using the public dataset MMSE-HR. To deal with the size of the MMSE-HR database, we employed the 5-fold cross-validation method for heart rate estimation, and the folds were split independently between subjects to demonstrate generalizability. Table. 1 presents the average heart rate estimation performance of facial and acral rPPG. The proposed method achieved MAE, RMSE, and r of 1.85, 4.51, and 0.948, respectively, and outperformed previous deep learning-based approaches. The Pearson correlation of predicted and actual heart rate based on facial rPPG per fold is shown in Fig. 5. This graphs represent a linear trend, indicating a significant similarity between the predicted and actual heart rate. Algorithms (Nowara et al. (2021), Jaiswal & Meenpal (2022), Ouzar et al. (2023)) with short input image sequences provide fast inference times but at the cost of higher error rates. Video transformer-based algorithms such as Yu et al. (2022) and Yu et al. (2023) show considerable performance in cross-dataset tests. However, they require high-performance computing equipment due to their extensive parameters and computational demands. Additionally, Fig. 6 shows the 2 predicted rPPG and their phase differences as the cardiac cycle increases. Comparing the 2 approaches in our proposed method in Table. 1, it indicates that the heart rate estimation error based on the facial rPPG is lower than the acral rPPG. These experimental findings suggest that forcing the rPPG phase leads to decreased model performance and emphasize the importance of inferring actual pulse waves from captured region.
@@
The proposed BBP-Net for blood pressure estimation also outperformed the existing methods, achieving MAE and RMSE of 10.19 and 13.01 for SBP and 7.09 and 8.86 for DBP, respectively. Table. 2 presents the results of the blood pressure estimation experiment on the MMSE-HR database, and Fig. 7 shows the results of the Bland-Altman plot. The Rong & Li (2021) and Schrumpf et al. (2021) algorithms extracted rPPG using a mathematical approach and estimated blood pressure through handcrafted features and deep learning-based models, respectively. These studies are similar to our approach in extracting rPPG as an intermediate step. However, our method improves performance by considering phase-shifted information for blood pressure estimation. Chen et al. (2023) proposed a blood pressure estimation model based on a 2 dimensional spatiotemporal map, a video preprocessing technique based on the YUV color space. However, they neglect pulse waves at the acral sites of the body, resulting in higher SBP and DBP error rates compared to the proposed method. The experimental results in Table. 2 represent that the temporal information in the pulse wave is meaningful for non-contact blood pressure estimation.
@@ 4.3 Quantitative results on V4V database
In this section, we conducted experiments on V4V database to compare the heart rate estimation performance of the proposed DRP-Net with existing algorithms and the experimental results are shown in Table. 3. The experiments are conducted on the training and test data, and the graph 6 in Fig. 5 shows a Pearson correlation plot of the predicted and actual heart rates. The metrics for DRP-Net heart rate estimation performance using V4V are the same as for the MMSE-HR database. The proposed method using face rPPG outperformed in MAE and r, with 3.83 and 0.75, respectively. Additionally, the proposed method with acral rPPG achieved MAE, RMSE, and r of 4.13, 10.14, and 0.72, respectively This result is a slightly higher error than facial rPPG, similar to the results on the MMSE-HR database. The performance comparison between Yu et al. (2019) and DRP-Net shows that the wider receptive fields through 3D atrous convolution is beneficial for heart rate estimation than the encoder-decoder structure based on 3D CNN. The performance comparison with Kim et al. (2022) shows contrasting MAE and RMSE results. It indicates that the outliers of heart rate predicted by the proposed method are larger, while the average error is smaller.
@@
The blood pressure estimation results of BBP-Net is shown in Table. 4. Our proposed algorithm achieved MAE and RMSE of 13.64, 16.78 and 9.40, 11.90 for SBP and DBP, respectively. The performance of the baseline algorithms (AlexNet, ResNet, LSTM) is reported in Hamoud et al. (2023). Blood pressure estimation performance of BBP-Net outperforms baseline algorithms such as AlexNet, ResNet, and LSTM, as well as Schrumpf et al. (2021) and the algorithm proposed by Hamoud et al. (2023). While the SBP and DBP estimation performance presented in the Hamoud et al. (2023) has the limitation that each performance is a combination of different models, the BBP-Net estimates both blood pressures simultaneously. As shown in the experimental results on the V4V database, the rPPG estimated from the 2 body parts and its phase-shifted information are effective in improving the estimation performance of blood pressure.
@@ 4.4 Ablation study
Table 5: Comparative results of mean absolute error in heart rate estimation task across various window sizes and folds.
@@
Table 6: Comparison of MAE in heart rate estimation with and without data augmentation.
Augmentation rPPG MAE (BPM)
X facial 2.149
@@
Table 7: Comparison of MAE in heart rate estimation with and without peak valley loss.
Peak valley loss rPPG MAE (BPM)
X facial 3.670
acral 4.178
O facial 1.783
acral 1.909
The window size of the input image sequence is an important heuristic parameter in terms of the computational power and the latency of model. Smaller window sizes have the advantage of lower latency in inference time, while the loss of temporal information can decrease the performance. Table. 5 presents the performance with different window size and fold, and it demonstrates that the average heart rate estimation performance increases as the window size increases. The reduced performance observed at a window size of 2 seconds is due to the dilation ratio of DRP-Net being relatively large compared to the sequence information. However, estimation of blood pressure requires long temporal information of rPPG, therefore we selected window size of 6 seconds and a dilation ratio of 4. Table. 6 indicates the results of heart rate estimation of DRP-Net with and without data augmentation techniques. The performance of all estimated rPPG improved after data augmentation. This result demonstrate that training with data from bradycardia and tachycardia via frame interpolation and down-sampling improves the generalization performance of the model. In addition, the experimental results of excluding the peak and valley guides from the loss function are demonstrated in Table. 7 to investigate the impact of $L_{peakvalley}$. The average MAE without $L_{peakvalley}$ increases about 2 BPM, which indicates that the amplitude information in the PPG contributes the improvement of heart rate estimation performance.
Table 8: Comparative results of MAE in blood pressure estimation for the input rPPG types.
@@
Table 9: Comparison of blood pressure estimation performance with and without the scaling sigmoid function.
Scaling sigmoid function SBP DBP
@@
The ablation study for the blood pressure estimation model consists of the types of input rPPG and the existence of the scaling sigmoid function. Table. 8 presents the estimation error of the SBP and DBP based on the facial or acral rPPG, and the estimation performance of both cases are lower than using dual rPPG. This results indicate that the phase-shifted information of pulse waves in different body parts contributes to improved estimation performance of SBP and DBP. Additionally, Table. 9 shows the performance of blood pressure estimation with and without the scaling sigmoid function. Without the scaling sigmoid function, the MAE for SBP and DBP was significantly increased to 9 mmHg and 3 mmHg, respectively. These experimental results suggest the importance of guiding the model towards a reasonable blood pressure range.
4.5 International standards and blood pressure estimation performance
Table 10: Comparison results of blood pressure estimation performance with BHS standards.
@@ 4.5 International standards and blood pressure estimation performance
Table 11: Comparison results of blood pressure estimation performance with AAMI standards.
@@
We compare the performance of our proposed blood pressure estimation algorithm with international standards, the British Hypertension Society (BHS)  (O’Brien et al., 1990) and The Association for the Advancement of Medical Instrumentation (AAMI)  (Stergiou et al., 2018). The BHS standard evaluates the absolute number of blood pressure estimation errors as a cumulative percentage, and Table. 10 presents the standards and blood pressure estimation performance. On the MMSE-HR database, the proposed method achieves grade B with 81 % DBP estimation performance for errors of 5 mmHg or less. Additionally, DBP estimation performance on the V4V database also reached grade C. The AAMI standard in Table. 11 indicates that blood pressure estimation performance meets the standard when the mean error (ME) and standard deviation error (SDE) are under 5 and 8 mmHg, respectively. In the MMSE-HR database, the estimation performance of SBP and DBP for ME is -0.13 mmHg and 1.58 mmHg, respectively, which meets the standard. The result on the V4V database shows that ME is 5.96 mmHg at SBP and -2.4 mmHg at DBP.
@@ 5 Conclusion
This paper proposed the deep learning framework based on facial images for contactless heart rate and blood pressure estimation. We assumed that phase differences exist between pulse waves due to inconsistencies in data measurement sites. Based on this assumption, we improved blood pressure estimation performance and demonstrated with extensive experiments. To estimate pulse waves of 2 physiological sites from facial images, we proposed the DRP-Net that focuses on spatiotemporal feature extraction. DRP-Net estimates facial rPPG using information in the time-global frequency domain to predict pulse waves from the face, and predicts acral rPPG that mimic the phase of PPG. As a result, the proposed model achieved the MAE of 1.78 and 1.91 BPM based on acral and facial rPPG in heart rate estimation task. Moreover, we proposed BBP-Net to estimate blood pressure based on the phase analysis and temporal features of the dual rPPG, applying the modified sigmoid function to bound the predicted blood pressure to the reasonable range. In the blood pressure estimation task, we achieved MAE of SBP and DBP of 10.19 mmHg and 7.09 mmHg, respectively. However, the public datasets used in this study have the limitation in that they do not include pulse wave data from the face, making it unable to verify the accuracy of the predicted facial rPPG phase. We were also constrained by the absence of other public datasets containing both facial images and blood pressure data, limiting experimentation in various domains. Inspired by these limitations, we have plan to collect real-world dataset which includes facial videos, PPG, and blood pressure with various data collection scenarios and develop sensor acquisition system in our future research. Additionally, we will conduct research on improving the performance of physiological measurement system using the collected real-world dataset.

### 06 Mar 2024 v2

@@ 4.1 Datasets
We conducted our experiment using the MMSE-HR (Multimodal Spontaneous Expression-Heart Rate) (Zhang et al., 2016b) and V4V (Vision for Vitals) database (Revanur et al., 2021) to demonstrate the heart rate and blood pressure estimation performance of the proposed method. The utilized datasets are sub-datasets derived from the MMSE database (BP4D+), which consists of synchronized facial image sequences and continuous ABP signals. As introduced in Section 3.1, the physiological information of pseudo PPG signals, heart rate, SBP, and DBP can be calculated from ABP signals. For physiological diversity in the dataset, subjects were conducted various tasks to arouse amusement, pain, anger, and disgust. RGB video was recorded at the resolution of 1040$\times$1392 with a frame rate of 25 FPS, and ABP signals were collected at the sampling rate of 1000 Hz. The MMSE-HR database released at the CVPR 2016 for The Heart Rate Estimation Algorithm Challenge includes 102 samples from 17 male and 23 female subjects. The average length of signals is 30 seconds. The V4V database released at the ICCV workshop 2021 for The 1st Vision for Vitals Challenge includes 1,358 data samples from 179 subjects split into 724 for training, 276 for validation, and 358 for test samples. The average length of signals is 40 seconds.
@@
Figure 5: Pearson correlation graph of predicted and actual heart rate
Figure 6: Predicted facial (blue), acral (green) rPPG and pseudo PPG (red) signals
@@ 4.2 Quantitative results on MMSE-HR database
In this section, we compare the heart rate estimation performance of the DRP-Net with previous methods using the public dataset MMSE-HR. To deal with the small size of the MMSE-HR database, we employed the 5-fold cross-validation method for heart rate estimation, and the folds were split independently between subjects to demonstrate generalizability. Table. 1 presents the average of fold-wise heart rate estimation performance of facial and acral rPPG signals. The proposed method achieved MAE, RMSE, and r of 1.78, 4.27, and 0.95, respectively, and outperformed previous deep learning-based approaches. The Pearson correlation of actual and predicted heart rate based on facial rPPG signals is shown in Fig. 5. These graphs show a linear trend, indicating a significant similarity between the predicted and actual heart rate. Algorithms (Nowara et al. (2021), Jaiswal & Meenpal (2022), Ouzar et al. (2023)) with short input image sequences provide fast inference times. However, the loss of temporal information leads to an increase in estimation error. Video transformer-based algorithms such as Yu et al. (2022) and Yu et al. (2023) demonstrate significant performance in cross-dataset tests. However, transformers require high-performance computing resources due to their extensive parameters and computational demands. Additionally, Fig. 6 shows the predicted facial and acral rPPG signals and their temporal discrepancy. Comparing the two approaches in our proposed method in Table. 1, the results indicate that the heart rate estimation error based on the facial rPPG signals is lower than the acral rPPG signals. These experimental findings suggest that forcing the phase of rPPG signal leads to decreased model performance and emphasize the importance of inferring actual pulse waves from a captured region.
@@
The proposed BBP-Net for blood pressure estimation also outperformed the existing methods, achieving MAE and RMSE of 10.19 and 13.01 for SBP and 7.09 and 8.86 for DBP, respectively. Table. 2 presents the results of the blood pressure estimation experiment on the MMSE-HR database. Fig. 7 shows the blood pressure estimation results of the Bland-Altman plots, and the solid line and dotted lines represents mean of difference and 95% limits of agreement, respectively. The algorithms proposed by Rong & Li (2021) and Schrumpf et al. (2021) extracted rPPG signals using a non-parametric approach and estimated blood pressure via handcrafted feature-based and deep learning-based models, respectively. These studies are similar to our approach in extracting rPPG signals as an intermediate step. However, the difference of our method lies in the utilization of phase-shifted rPPG signals derived from a parametric model. Chen et al. (2023) proposed a blood pressure estimation model that utilizes a two-dimensional spatiotemporal map, a video preprocessing technique based on various color spaces. However, they neglect to consider temporal discrepancy from the cardiac cycle, leading to increased errors in estimating SBP and DBP. The experimental results in Table. 2 represent that the temporal discrepancy in the rPPG signals is meaningful for non-contact blood pressure estimation.
@@ 4.3 Quantitative results on V4V database
In this section, we present experiments conducted on the V4V database to compare the heart rate estimation performance of the proposed DRP-Net with existing algorithms, as shown in Table 3. Facial rPPG signals derived from the proposed method outperformed others in terms of MAE and r, with values of 3.83 and 0.75, respectively. The sixth graph in Fig. 5 shows a Pearson correlation plot of the predicted and actual heart rates. Additionally, the proposed method with acral rPPG signals achieved MAE, RMSE, and r of 4.13, 10.14, and 0.72, respectively. This result indicates a slightly higher error compared to the results of facial rPPG signals, consistent with findings from the MMSE-HR database. The performance comparison between Yu et al. (2019) and DRP-Net indicates that utilizing wider receptive fields through atrous convolution is more advantageous for learning spatiotemporal features than the encoder-decoder structure based on three dimensional CNN. The performance comparison with Kim et al. (2022) shows contrasting MAE and RMSE results. This result indicates that the error of outliers of proposed method is large.
The blood pressure estimation results of BBP-Net is shown in Table. 4. Our proposed algorithm achieved MAE, RMSE of 13.64, 16.78 and 9.40, 11.90 for SBP and DBP, respectively. Blood pressure estimation performance of BBP-Net outperforms baseline algorithms such as AlexNet, ResNet, and LSTM, as well as Schrumpf et al. (2021) and the algorithms proposed by Hamoud et al. (2023). The performance of the baseline algorithms (AlexNet, ResNet, LSTM) is reported in Hamoud et al. (2023). While the SBP and DBP estimation performance presented in the Hamoud et al. (2023) has the limitation that each performance is a combination of different models, the BBP-Net estimates both of blood pressures simultaneously. As shown in the experimental results on the V4V database, the rPPG signals estimated from the two physiological sites and its phase-shifted information are effective in improving the estimation performance of blood pressure.
@@ 4.4 Ablation study
Table 5: Comparative results of heart rate estimation performance across various window lengths and folds.
@@
Table 6: Comparative results of heart rate estimation performance with data augmentation and $L_{pv}$.
Augmentation $L_{pv}$ rPPG MAE (BPM)
facial 2.149
@@
Table 7: Comparative results results of blood pressure estimation performance for the type of rPPG.
@@
Table 8: Comparative results of blood pressure estimation performance with and without the scaled sigmoid function.
Scaled sigmoid function SBP DBP
@@
The ablation study on heart rate estimation is conducted to investigate the impact of the proposed methods. The window length of the input image sequence is an important heuristic parameter in terms of the computational power and the latency of the model. Smaller window length have the advantage of lower latency in inference time, however, the loss of temporal information can decrease performance. Table. 5 presents the performance with different window length, and demonstrates that the heart rate estimation performance increases as the window length increases. The lowest performance observed at a window length of 2 seconds is due to the dilation ratio of DRP-Net being relatively large compared to the sequence length. However, estimation of blood pressure requires long-term information of physiological signals, therefore we selected a window length of 6 seconds and a dilation ratio of 4. Table. 6 indicates the results of heart rate estimation of DRP-Net with and without data augmentation techniques and $L_{pv}$. The performance of rPPG estimation improved after data augmentation. This result demonstrate that training with data from bradycardia and tachycardia via frame interpolation and down-sampling improves the generalizability. Additionally, the experimental results of excluding the peak and valley guides from the loss function are also demonstrated in Table. 6 to investigate the impact of $L_{pv}$. The average MAE with $L_{pv}$ decreases by about 2 BPM, which indicates that the guidance of scale significantly affects the estimation of accurate heart rate.
The ablation study for BBP-Net focuses on the input types of rPPG and the utilization of the scaled sigmoid function. Table 7 presents the estimation errors for blood pressure from different input types of rPPG signals and shows the superior performance of phase-shifted rPPG. This results indicate that the temporal discrepancy of pulse waves in different physiological sites contributes to improved estimation performance of blood pressure. Table. 8 shows the performance of blood pressure estimation with and without the scaled sigmoid function of BBP-Net. Without the scaled sigmoid function, the MAE for SBP and DBP increased by 9 mmHg and 3 mmHg, respectively. These experimental results suggest the importance of guiding the model towards a reasonable blood pressure range.
4.5 International standards
Table 9: Comparison of blood pressure estimation performance with BHS standards.
@@ 4.5 International standards
Table 10: Comparison of blood pressure estimation performance with AAMI standards.
@@
We compare the performance of proposed algorithm with international standards, the British Hypertension Society (BHS)  (O’Brien et al., 1990) and The Association for the Advancement of Medical Instrumentation (AAMI)  (Stergiou et al., 2018). The BHS standard evaluates the absolute number of blood pressure estimation errors as a cumulative percentage, as shown in Table. 9. On the MMSE-HR database, the proposed method achieves grade B with 81 % DBP estimation performance for errors of 5 mmHg or less. Additionally, DBP estimation performance on the V4V database also achieved grade C. The AAMI standard in Table. 10 indicates that blood pressure estimation performance meets the standard when the mean error (ME) and standard deviation error (SDE) are under 5 and 8 mmHg, respectively. In the MMSE-HR database, the blood pressure estimation performances in terms of ME are -0.13 mmHg and 1.58 mmHg, respectively, which meet the standard.
@@ 5 Conclusion
This paper proposes a deep learning framework based on facial images for contactless heart rate and blood pressure estimation. We assumed temporal discrepancy between pulse waves due to potential inconsistencies in data measurement sites. To estimate phase-shifted rPPG signals, we propose DRP-Net which learns spatiotemporal features from facial images via frequency and time domain loss functions. Additionally, frame sampling based augmentation is proposed to augment bradycardia and tachycardia data. Moreover, we propose BBP-Net to estimate blood pressure from phase-shifted rPPG signals and its temporal discrepancy. In the experimental results, we demonstrated the improvement in heart rate and blood pressure estimation performance by previous methods and ablation studies. However, we found limitations in most public datasets, which exclude facial PPG signals or blood pressure data. Inspired by founded limitations, we are planning to collect a real-world dataset with various scenarios to verify phase of facial rPPG signals and the performance of the proposed deep learning framework. Finally, we hope that our experimental findings can be actively utilized in further research for robust remote physiological measurement.
Acknowledgements
Collate acknowledgements in a separate section at the end of the article before the references and do not, therefore, include them on the title page, as a footnote to the title or otherwise. List here those individuals who provided help during the research (e.g., providing language help, writing assistance or proof reading the article, etc.).